i need help with this second question two, theyre similar.


X represents the Grade Point Average (GPA) of students at a private college which follows normal distribution with mean of 2.8 and standard deviation of 0.52. If top 5% of the students will receive special awards, what is the minimum GPA required to qualify for the award?








1
Expert's answer
2021-09-02T12:35:20-0400

"\mu=2.8 \\

\sigma=0.52 \\

P(X<x) = 1-0.05 = 0.95 \\

P(Z< \frac{x -2.8}{0.52}) = 0.95"

From table of Z Score

"\frac{x -2.8}{0.52} = 1.65 \\

x -2.8 = 0.52 \times 1.65 \\

x = 2.8 + 0.858 \\

x = 3.658"

Answer: The minimum GPA required to qualify for the award is 3.66.